# A Unified Empirical Risk Minimization Framework for Flexible N-Tuples Weak Supervision

### v1

@@ Abstract
Index Terms:
Weakly-supervised learning, N-tuples learning, pointwise unlabeled data , unbiased risk estimator.
@@ I Introduction
In this work, we propose a unified N-tuple weakly-supervised learning framework that subsumes existing N-tuple methods and extends them with greater flexibility. We begin by considering $\bar{\mathcal{Y}}$ as the full label space consisting of all $2^{N}$ possible label configurations for an N-tuple of binary instances, where each instance is labeled as either positive ($+1$) or negative ($-1$). For any given weak supervision scenario, we then define a subset $\mathcal{Y}^{\text{sub}}\subseteq\bar{\mathcal{Y}}$ of these assignments that satisfy the task’s constraints. For example, the NT-Comp scenario corresponds to those assignments where the probabilities of being positive are strictly decreasing across the tuple, i.e., $\mathbb{P}(y_{1}=+1)>\mathbb{P}(y_{2}=+1)>\dots>\mathbb{P}(y_{N}=+1)$, while NSU corresponds to assignments where all labels in the tuple are identical, i.e., $\forall i,j\in\{1,2,\dots,N\},\ y_{i}=y_{j}$. We further introduce two more general scenarios that naturally combine N-tuple weak supervision with pointwise unlabeled data. Specifically, the mixed-class N-tuples and pointwise unlabeled learning (MNU) allows tuples where not all instances share the same label, i.e., $\exists i,j\in\{1,2,\dots,N\},\ y_{i}\neq y_{j}$; and the not-all-negative N-tuples and pointwise unlabeled learning ($\bm{N_{pos}U}$) ensures at least one positive instance, i.e., $\exists i\in\{1,2,\dots,N\},\ y_{i}=+1$. These settings correspond to practical tasks such as academic performance ranking (NT-Comp), batch quality inspection (NSU), general image classification (MNU), and fraud detection ($N_{pos}U$). By selecting different subsets of the label space, our framework unifies these diverse supervision forms and eliminates the need for task-specific model designs. The specific label constraints corresponding to each weakly supervised scenario are summarized in Table II.
@@ II Preliminaries
Supervised classification is a traditional learning paradigm that trains classifiers using precisely labeled examples. Given a dataset with precisely labeled instances, let $\mathcal{X}\subset\mathbb{R}^{d}$ denote the feature space consisting of both positive and negative examples. $\mathcal{Y}=\{-1,1\}$ indicates the label space of $\mathcal{X}$, with $y=1$ denoting a positive instance and $y=-1$ a negative one. The positive dataset $\mathcal{X}_{p}$ is independently drawn from the marginal distribution $p_{+}(\textbf{x})=p(\textbf{x}\mid y=+1)$. Similarly, the negative sample set $\mathcal{X}_{n}$ is independently sampled from the marginal distribution $p_{-}(\textbf{x})=p(\textbf{x}\mid y=-1)$.
@@ III-C Estimation Error Bound
4:     for $j=1,2,$ number of batch_size do
@@ IV-A Case 1: N-tuples comparisons and unlabeled learning ($N_{\text{Comp}}U$)
$\begin{split}C_{1j}&=\dfrac{\alpha_{j}\gamma_{3}-\beta_{j}\gamma_{2}}{\gamma_{1}\gamma_{3}-\gamma_{2}^{2}},\quad C_{2j}=\dfrac{-\alpha_{j}\gamma_{2}+\beta_{j}\gamma_{1}}{\gamma_{1}\gamma_{3}-\gamma_{2}^{2}},\\ D_{1}&=\frac{\tau_{+}\gamma_{3}-\tau_{-}\gamma_{2}}{\gamma_{1}\gamma_{3}-\gamma_{2}^{2}},\quad D_{2}=\frac{-\tau_{+}\gamma_{2}+\tau_{-}\gamma_{1}}{\gamma_{1}\gamma_{3}-\gamma_{2}^{2}}.\\ \end{split}$ (23)
@@ IV-B Case 2: N-tuples similarities and unlabeled learning
$\begin{split}&\widehat{R}_{s}(g)=\frac{\tau_{+}^{N}+\tau_{-}^{N}}{(\tau_{+}^{N-1}-\tau_{-}^{N-1})n_{s}^{n}N}\sum_{i=1}^{n_{s}^{n}N}[\ell(g(\textbf{x}),+1)-\ell(g(\textbf{x}),-1)]\\ &+\frac{1}{n_{u}}\sum_{i=1}^{n_{u}}[-\frac{\tau_{-}^{2}}{2\tau_{+}-1}\ell(g(\textbf{x}),+1)+\frac{\tau_{+}^{2}}{2\tau_{+}-1}\ell(g(\textbf{x}),-1)],\\ \end{split}$ (27)

### v2

@@ I Introduction
In this work, we propose a unified N-tuple weakly-supervised learning framework that subsumes existing N-tuple methods and extends them with greater flexibility. We begin by considering $\bar{\mathcal{Y}}$ as the full label space consisting of all $2^{N}$ possible label configurations for an N-tuple of binary instances, where each instance is labeled as either positive (+1+1) or negative (1-1). For any given weak supervision scenario, we then define a subset $\mathcal{Y}^{\text{sub}}\subseteq\bar{\mathcal{Y}}$ of these assignments that satisfy the task’s constraints. For example, the NT-Comp scenario corresponds to those assignments where the probabilities of being positive are strictly decreasing across the tuple, i.e., $\mathbb{P}(y_{1}=+1)>\mathbb{P}(y_{2}=+1)>\dots>\mathbb{P}(y_{N}=+1)$, while NSU corresponds to assignments where all labels in the tuple are identical, i.e., $\forall i,j\in\{1,2,\dots,N\},\ y_{i}=y_{j}$. We further introduce two more general scenarios that naturally combine N-tuple weak supervision with pointwise unlabeled data. Specifically, the mixed-class N-tuples and pointwise unlabeled learning (MNU) allows tuples where not all instances share the same label, i.e., $\exists i,j\in\{1,2,\dots,N\},\ y_{i}\neq y_{j}$; and the not-all-negative N-tuples and pointwise unlabeled learning ($\bm{N_{pos}U}$) ensures at least one positive instance, i.e., $\exists i\in\{1,2,\dots,N\},\ y_{i}=+1$. These settings correspond to practical tasks such as academic performance ranking (NT-Comp), batch quality inspection (NSU), general image classification (MNU), and fraud detection ($N_{pos}U$). By selecting different subsets of the label space, our framework unifies these diverse supervision forms and eliminates the need for task-specific model designs. The specific label constraints corresponding to each weakly supervised scenario are summarized in Table II.
@@ II Preliminaries
Supervised classification is a traditional learning paradigm that trains classifiers using precisely labeled examples. Given a dataset with precisely labeled instances, let $\mathcal{X}\subset\mathbb{R}^{d}$ denote the feature space consisting of both positive and negative examples. $\mathcal{Y}=\{-1,1\}$ indicates the label space of $\mathcal{X}$, with $y=1$ denoting a positive instance and y=1y=-1 a negative one. The positive dataset $\mathcal{X}_{p}$ is independently drawn from the marginal distribution $p_{+}(\textbf{x})=p(\textbf{x}\mid y=+1)$. Similarly, the negative sample set $\mathcal{X}_{n}$ is independently sampled from the marginal distribution $p_{-}(\textbf{x})=p(\textbf{x}\mid y=-1)$.
@@ III-C Estimation Error Bound
4:  for $j=1,2,$ … number of batch_size do
@@ IV-A Case 1: N-tuples comparisons and unlabeled learning ($N_{\text{Comp}}U$)
[…]
@@ IV-B Case 2: N-tuples similarities and unlabeled learning
[…]